# Yong Ji1, Junye Li1 and Rui Yang∗2,3,4

### v2

@@ 3.2 Bowen metric mean dimension of saturated sets
We prove Theorems 1.2 and 1.3 in this subsection.
@@ 3.2.1 Upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$
The following lemma gives an upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$, which was already given in [Y23, Lemma 3.3,(3)]
@@
Let $(X,f)$ be a t.d.s.. Let $K\subset M_{f}(X)$ be a nonempty closed connected subset. Then for any $\epsilon>0$, we have
@@
Thus, by the Lemma 3.6 we obtain an upper bound of the Bowen metric mean dimension of saturated set:
@@ 3.2.2 Lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
The proof of the lower bound is similar to that of in subsection 3.1.2. So we only give a sketch for the estimation of lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$.
We shall show that for any small enough $\epsilon^{\ast}>0$, it holds that
@@
Put $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$ and endow
$\mathcal{A}=\{(k,0):k\geq 0\}\cup\mathcal{A}^{\prime}$
with the lexicographic order again. Using the fixed notation in subsection 3.1.2, we have the data $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ and $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$.
@@
where $n_{(k,i)}:=n_{(k,i)}^{\prime}+(p_{k,i}-1)m_{(k,i)}$.
$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$ which are $(\lambda_{k,i,l}n_{(k,i)}^{\prime},5\epsilon^{\ast})$-separated and
@@
Then $\mathcal{C}_{(k,i)}$ is a $(n_{(k,i)},4\epsilon^{\ast})$-separated set with
@@
Set $\mathcal{S}_{(k,0)}=\mathcal{C}_{(k,0)}$, $c_{k,0}=n_{(k,0)}$ for $k\geq 0$. For $(k,i)\in\mathcal{A}^{\prime}$, and any tuple of points $(y_{k,i,1},y_{(k,i,2)},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists a point $z=z(y_{k,i,1},y_{(k,i,2)},\ldots,y_{k,i,N_{(k,i)}})$ that $\epsilon_{(k,i)}$-shadows the pieces of orbits
@@
with the same gap $m_{(k,i)}$. Denote by $\mathcal{S}_{(k,i)}$ the set of such shadowing points, which is $(c_{k,i},3\epsilon^{\ast})$-separated set, where $c_{k,i}=n_{(k,i)}N_{(k,i)}+(N_{(k,i)}-1)m_{(k,i)}$.
@@
Without loss of generality, we can assume that $\rho_{(k,j)}\to\rho$. Furthermore, one can check that $\rho(\mathbb{F})=1$, and for every $n$ large enough and any ball $B_{n}(z,\frac{\epsilon^{\ast}}{2})$ intersecting with $\mathbb{F}$, it holds that
@@
Then by the Lemma 2.1, we have
@@
$\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{B}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}\xi<\epsilon}h_{\mu}(f,\xi).$ (3.15)
@@
On the other hand, we may assume that $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)$ is finite; otherwise there is nothing left to prove. Let $\gamma>0$. Then according to Theorem 1.2 we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence of $\{\mu_{k}\}$ in $K$ such that for any $k$,
@@
Assume that $\mu_{k}\rightarrow\mu$. Then $\mu\in M_{f}(X)$ since $K$ is closed. This shows
$\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
@@
Using formula (3.2), the inequality
$\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}\leq\overline{\rm mdim}_{M}^{P}(f,G_{K},d).$
is clear. On the other hand, let $\gamma>0$ be sufficiently small and put $M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}$.Then we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence of $\{\mu_{k}\}$ in $K$ such that for sufficiently large $k$,
@@
Again, we assume that $\mu_{k}\rightarrow\mu\in K$. This shows
$M_{\gamma}\leq\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
@@ 3.3 Upper capacity metric mean dimension of saturated sets
We prove the Theorem 1.4 in this subsection.
The inequality
$\overline{\rm mdim}_{M}^{UC}\left(f,G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},d\right)\leq\overline{\rm mdim}_{M}^{UC}\left(f,G_{K},d\right)\leq\overline{\rm mdim}_{M}(f,X,d)$
@@
$\overline{\rm mdim}_{M}(f,X,d)\leq\overline{\rm mdim}_{M}^{UC}\left(f,G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},d\right).$
Once we prove that for $\epsilon^{\ast}>0$, $\zeta>0$ small enough and $\delta\in(0,1)$, it holds that
$\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta\leq h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast}),$
then by Theorem 2.4, the proof of Theorem 1.4 will be done. Since the proof is similar to the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$ in subsection 3.1.2, we give a sketch for it.
We write $H=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)>0$. Take $\mu^{\ast}\in M_{f}^{e}(X)$ such that
@@
Then for large enough $\mathcal{N}\in\mathbb{N}$, there exists a $(\mathcal{N},3\epsilon^{\ast})$-separated subset $S_{\mathcal{N}}$ with
@@
Since $K$ is a connected closed subset, a standard argument shows that there exists a family $\{\mu_{k,i}:k\geq 1,1\leq i\leq Q_{k}\}\subset K$ of measures in $K$ such that for all $k\geq 1$, $1\leq i<Q_{k}$,
@@
Consider the sets $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$, and
$\mathcal{A}=\{(k,0):k\in\mathbb{Z}\ \text{and }k\geq-1\}\cup\mathcal{A}^{\prime}$
with the lexicographic order. The sequences $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ are constructed similarly as in subsection 3.1.2.
@@
Recall that $M_{(0,0)}=1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}$. Hence for $y$, $y^{\prime}\in\mathbb{F}$ with $y\in\overline{B}_{M_{(0,0)}}(x,2\epsilon_{(0,0)})$, and $y^{\prime}\in\overline{B}_{M_{(0,0)}}(x^{\prime},2\epsilon_{(0,0)})$ one has
@@
Hence for any large enough $\mathcal{N}$, there is a $(1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N},\epsilon^{\ast})$-separated subset in $\mathbb{F}$ with cardinality #$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)]$. Thus
@@
4 Applications

### v3

@@ 3.2 Bowen metric mean dimension of saturated sets
In this subsection, we prove Theorems 1.2 and 1.3.
@@ 3.2.1 Upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$
The following lemma gives an upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$, which was already obtained in [Y23, Lemma 3.3,(3)].
@@
Let $(X,f)$ be a t.d.s., and let $K\subset M_{f}(X)$ be a nonempty closed subset.
  • (1)
    For any $\epsilon>0$, we have
    $h_{top}^{B}(f,\{x\in X:V_{f}(x)\cap K\not=\emptyset\},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
  • (2)
    For any $\epsilon>0$, we have
@@
Thus, by the Lemma 3.6 the upper bound of the Bowen metric mean dimension of saturated set is given by:
@@ 3.2.2 Lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
The proof for the lower bound is similar to that in subsection 3.1.2. So we only give a sketch for the estimation of the lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$.
We show that for any small enough $\epsilon^{\ast}>0$, it holds that
@@
Put $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$, $\mathcal{A}=\{(k,0):k\geq 0\}\cup\mathcal{A}^{\prime}$, and endow these sets with the lexicographic order again. Using the fixed notation in subsection 3.1.2, we have the data $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ and $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$.
@@
where $n_{(k,i)}:=n_{(k,i)}^{\prime}+(p_{k,i}-1)m_{(k,i)}$;
$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$, which are $(\lambda_{k,i,l}n_{(k,i)}^{\prime},5\epsilon^{\ast})$-separated and
@@
with gap $m_{(k,i)}$. Then $\mathcal{C}_{(k,i)}$ is a $(n_{(k,i)},4\epsilon^{\ast})$-separated set with
@@
Set $\mathcal{S}_{(k,0)}=\mathcal{C}_{(k,0)}$, $c_{k,0}=n_{(k,0)}$ for $k\geq 0$. For $(k,i)\in\mathcal{A}^{\prime}$, and any tuple of points $(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists a point $z=z(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})$ that $\epsilon_{(k,i)}$-shadows the pieces of orbits
@@
with the same gap $m_{(k,i)}$. Denote by $\mathcal{S}_{(k,i)}$ the set of such shadowing points. It is a $(c_{k,i},3\epsilon^{\ast})$-separated set, where $c_{k,i}=n_{(k,i)}N_{(k,i)}+(N_{(k,i)}-1)m_{(k,i)}$.
@@
Without loss of generality, we can assume that $\rho_{(k,j)}\to\rho$. Furthermore, one can check that $\rho(\mathbb{F})=1$, and for every $n$ large enough and any ball $B_{n}(z,\frac{\epsilon^{\ast}}{2})$ intersecting with the fractal set $\mathbb{F}$, it holds that
@@
Then, by the Lemma 2.1, we have
@@
$\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{B}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}\xi<\epsilon}h_{\mu}(f,\xi).$
@@
On the other hand, we may assume that $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)$ is finite; otherwise there is nothing left to prove. Let $\gamma>0$. Then, by Theorem 1.2, we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for any $k$,
@@
Assume that $\mu_{k}\rightarrow\mu$ as $k\to\infty$. Then $\mu\in K$ since $K$ is closed. This shows that
$\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
@@
Next we show the second equality:
Using formula (3.3), the inequality $\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}\leq\overline{\rm mdim}_{M}^{P}(f,G_{K},d)$ is clear. On the other hand, let $\gamma>0$ be sufficiently small and put
$M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}.$
Then we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for sufficiently large $k$,
@@
Again, we assume that $\mu_{k}\rightarrow\mu\in K$. This shows that
$M_{\gamma}\leq\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}.$
@@ 3.3 Upper capacity metric mean dimension of saturated sets
In this subsection, we prove the Theorem 1.4.
By Proposition 2.1, the inequality
$\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d)\leq\overline{\rm mdim}_{M}^{UC}(f,G_{K},d)\leq\overline{\rm mdim}_{M}(f,X,d)$
@@
$\overline{\rm mdim}_{M}(f,X,d)\leq\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d).$
In light of Theorem 2.4, we need to show that for $\epsilon^{\ast}>0$, $\zeta>0$ sufficiently small and $\delta\in(0,1)$, one has
$\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta\leq h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast}).$ (3.19)
However, we will give a sketch for the proof of (3.19) since it is essentially analogous to that of the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$ presented in subsection 3.1.2.
Write $H=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)$ and assume that $H>0$. Take $\mu^{\ast}\in M_{f}^{e}(X)$ such that
@@
Then, for sufficiently large $\mathcal{N}\in\mathbb{N}$, there exists a $(\mathcal{N},3\epsilon^{\ast})$-separated subset $S_{\mathcal{N}}$ with
@@
Since $K$ is a connected closed subset, a standard argument shows that there exists a family $\{\mu_{k,i}:k\geq 1,1\leq i\leq Q_{k}\}$ of measures in $K$ such that for all $k\geq 1$, $1\leq i<Q_{k}$,
@@
Consider the set $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$ and endow $\mathcal{A}=\{(k,0):k\in\mathbb{Z}\ \text{and }k\geq-1\}\cup\mathcal{A}^{\prime}$ with the lexicographic order. Similarly, the sequences $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ are constructed as in subsection 3.1.2.
@@
Recall $M_{(0,0)}=1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}$. Let $y$, $y^{\prime}\in\mathbb{F}$ with $y\in\overline{B}_{M_{(0,0)}}(x,2\epsilon_{(0,0)})$ and $y^{\prime}\in\overline{B}_{M_{(0,0)}}(x^{\prime},2\epsilon_{(0,0)})$. Then one has
@@
Hence, for sufficiently large $\mathcal{N}$, there is a $(1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N},\epsilon^{\ast})$-separated subset in $\mathbb{F}$ with cardinality #$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)]$. Thus
@@
This implies the inequality (3.19).
3.4 An extension to metric mean dimensions of sup sets and sub sets
In this section, we consider two variants of saturated sets, namely sup sets and sub sets, and establish variational principles for metric mean dimensions of sup sets and sub sets. The main results are presented in Theorems 3.7, 3.8 and 3.9.
In general, given a point $x\in X$, the limit $\lim_{n\to\infty}\mathcal{E}_{n}(x)$ may not exist in the sense of weak-topology. The set of such points is denoted by $I(f)$, known as the historic set [^1] This terminology is due to Rulle’s original paper [Rue01]. of $f$. By Birkhoff ergodic theorem, although the historic set is usually regarded as negligible in the measure-theoretic context, its topological dynamics can be quite rich. Especially for systems with specification-like properties, it can even have full topological entropy [CTS05] and full metric mean dimension [LV21, LL24]. To understand the underlying geometry structure of $I(f)$, we define the following sets based on the framework introduced and further developed by Olsen [O02, O03, O06, O08] and Olsen and Winter [OW03]. More precisely, given a nonempty subset $C\subset M_{f}(X)$, we define the sup set, equ set [^2] In this context, we do not take into account the deformation of $\mathcal{E}_{n}(x)$ via a continuous affine map from $M_{f}(X)$ to a topological vector space $Y$. Hence, we prefer to use the term “saturated set” rather than “equ set”. and sub set of $C$ as
${}^{C}G:$ =$=\{x\in X:V_{f}(x)\supset C\},$
$G_{C}:$ =$=\{x\in X:V_{f}(x)=C\},$
$G^{C}:$ =$=\{x\in X:V_{f}(x)\subset C\},$
respectively.
Note that the equ set is precisely the saturated set associated with $C$ defined previously. More generally, by using two nonempty subsets $C_{1},C_{2}\subset M_{f}(X)$, we define the set
$G(C_{1},C_{2}):=\{x\in X:C_{1}\subset V_{f}(x)\subset C_{2}\},$
which covers the three different notions above in an unified manner.
Now, we are ready to establish variational principles for the three given types of level sets $G^{C},^{C}G,G(C_{1},C_{2})$ by means of the Bowen metric mean dimension.
Theorem 3.7.
Let $(X,d,f)$ be a t.d.s. with specification property. Then for any nonempty subset $C\subseteq M_{f}(X)$,
$\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
We claim that for every $\epsilon>0$,
$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)=\inf_{\mu\in\overline{{\rm Conv}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$ (3.20)
where $\overline{{\rm Conv}}(C)$ denotes the convex closure of set $C$.
Indeed, since $C\subset\overline{{\rm Conv}}(C)$, it suffices to show
$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\inf_{\mu\in\overline{{\rm Conv}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Let $\mu\in\overline{{\rm Conv}}(C)$. We can choose $\mu_{n}=\sum_{j=1}^{k_{n}}\lambda_{n}^{(j)}v_{n}^{(j)}$ with $v_{n}^{(j)}\in C$, $\lambda_{n}^{(j)}>0$ and $\sum_{j=o}^{k_{n}}\lambda_{n}^{(j)}=1$ such that $\mu_{n}\rightarrow\mu$ as $n\to\infty$. The upper semi-continuity of $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ on $M_{f}(X)$ (cf. [FW16, Lemma 2.3, (3)]) implies that
$\limsup_{n\to\infty}~{}{\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu_{n}}(f,\xi)}\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Combining this and $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ is concave in $\mu$, we can deduce that
$\inf_{\nu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\nu}(f,\xi)<\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)+\gamma$
holds for any for $\gamma>0$. This yields the desired inequality and hence (3.20) holds.
Notice that $G_{\overline{{\rm Conv}}(C)}\subset^{C}G$, we have
$\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm Conv}}(C)},d\right)$
[…]
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Conversely, for any $\mu\in C$ and $\epsilon>0$, by Lemma 3.6 one has
$h_{top}^{B}(f,^{C}G,\epsilon)\leq h_{top}^{B}(f,\{x\in X:\mu\in V_{f}(x)\},\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
This shows $h_{top}^{B}(f,^{C}G,\epsilon)\leq\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ for every $\epsilon>0$. ∎
Theorem 3.8.
Let $(X,d,f)$ be a t.d.s. with specification property. Then for any nonempty closed subset $C\subseteq M_{f}(X)$, we have
$\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
For every $\epsilon>0$, by Lemma 3.6 again we have
$h_{top}^{B}(f,G^{C},\epsilon)\leq h_{top}^{B}(f,\{x\in X:V_{f}(x)\cap{C}\not=\emptyset\},\epsilon)\leq\sup_{\mu\in{C}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
By (3.17), there exists a constant $c>0$ such that
$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu},\epsilon)$
holds for all $\mu\in C$ and any sufficiently small $\epsilon>0$. Combining with the fact that $G_{\mu}\subset G^{C}$ holds for each $\mu\in C$, one gets
$\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G^{C},\epsilon).$
Theorem 3.9.
Let $(X,d,f)$ be a t.d.s. with specification property. Suppose that $C_{1}\subseteq M_{f}(X)$ is a nonempty set satisfying that $\overline{{\rm Conv}}(C_{1})$ is a connected component of $C_{2}\subset M_{f}(X)$. Then
$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
Since $C_{1}\subset\overline{{\rm Conv}}(C_{1})\subset C_{2}$, by Theorem 1.2 we have
$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm Conv}}(C_{1})},d\right)$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in\overline{{\rm Conv}}(C_{1})}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Furthermore, we have $\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)\leq\overline{\rm mdim}_{M}^{B}\left(f,^{C_{1}}G,d\right)$. Using Theorem 3.7, we get
$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
This completes the proof. ∎
The Theorems 3.7, 3.8 and 3.9 imply the following corollaries, which allows us to compare the metric mean dimensions of sup sets and sub sets with the saturated sets.
Corollary 3.10.
Let $(X,d,f)$ be a t.d.s. with specification property.
  • (1)
    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)=\overline{\rm mdim}_{M}^{B}\left(f,G_{C},d\right).$
  • (2)
    If $C$ is a nonempty compact convex set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right).$
Proof.
(1) follows by Theorems 1.2 and 3.7; (2) is due to Theorems 1.1 and 3.8. ∎
Corollary 3.11.
Let $(X,d,f)$ be a t.d.s. with specification property.
  • (1)
    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{B}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)\leq\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right).$
  • (2)
    If $C$ is a nonempty compact convex set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right)\leq\overline{\rm mdim}_{M}^{P}\left(f,^{C}G,d\right).$
  • (3)
    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
    $\overline{\rm mdim}_{M}^{UC}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,G^{C},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,^{C}G,d\right).$
Proof.
$(1)$ is clear. For $(2)$, it follows from Theorem 1.1 that
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)=\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)\leq\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right).$
On the other hand, one has $h_{top}^{P}(f,G^{C},\epsilon)\leq\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ by Lemma 3.2. Hence, we have $\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right)$.
$(3)$ follows from Theorem 1.4 and the relations $G_{C}\subset G^{C}$ and $G_{C}\subset^{C}G$. ∎
4 Some applications of main results